OpenStudy (anonymous):

(2k + 7-5^3)(3k - 10+6^5)

OpenStudy (anonymous):

5k-70+11^5 is this right?

OpenStudy (anonymous):

Foil the problem

OpenStudy (anonymous):

(2k+3k) Front =5k (2k+6^5) Outside =7^5 (-5^3+3k) Inside = -2^3k (-5^3+6^5) Last. = 1^8 yea?

OpenStudy (anonymous):

the problem is multiplication, not addition

OpenStudy (anonymous):

its addition

OpenStudy (anonymous):

i don't see the addition on the problem

OpenStudy (anonymous):